Trip 1-: A boat goes upstream 30 km and downstream 44 km. The total time for this journey is 10hours. Trip 2-: The same boat also goes upstream 40 km and downstream 55 km . The total time for this trip is 13 hours. Find the speed of stream and boat. Answer-3 and 8 km/hr please give the solution

Trip 1-:
A boat goes upstream 30 km and downstream 44 km. The total time for this journey is 10hours.
Trip 2-:
The same boat also goes upstream 40 km and downstream 55 km . The total time for this trip is 13 hours.
Find the speed of stream and boat.
Answer-3 and 8 km/hr
please give the solution

Grade:11

3 Answers

Bhuvaneswari Singh
10 years ago
hi the equation is 30/x+y + 44/x+y = 10 similar for d next check rd sharma fr full solution.
Sunil Kumar FP
10 years ago
let the speed of the boat be v and the speed of the stream be s
according to the problem
44/v+s + 30/v-s =10
55/v+s + 40/v-s =13
let v+s=x v-s=y
44/x +30/y =10
55/x + 40/y =13
solving for x and y we get
x=11
y=5
now x=v+s
y=v-s
again solving v=8 a=3

thanks and regards
sunil kr
askIItian faculty

Sunil Raikwar
10 years ago
Dear Student, There is a technical issue so please re-post the question again separately. askIITian Faculty